OpenStudy (anonymous):

7x - 3y = 4 2x - 4y = 1 The solution to the system of equations is _____?

OpenStudy (anonymous):

Well first make one variable cancel out you can do this by multiplying the top by -2 and bottom by 7 for x's to cancel OR multiplying top by 4 and bottom by -3 to make y's cancel out.

OpenStudy (anonymous):

So if I multiplied -2 to the top one I would have -14x+6y=-8?

OpenStudy (anonymous):

yes and the bottom youd get 14x-28y=7 if you multipled by 7 and the x's cancel out and your left with -22y=-1 and if you divide by -22 you get y=1/22

OpenStudy (anonymous):

Okay.